# Observable Propagation: A Data-Efficient Approach to Uncover Feature Vectors in Transformers

### v1

@@ 4 Experiments
Armed with our “observable propagation” toolkit for obtaining and analyzing feature vectors, we now turn our attention to the problem of gender bias in LLMs in order to determine the extent to which these tools can be used to diagnose the causes of unwanted behavior.
@@ 4.1 Gendered pronouns prediction
We first consider the related question of understanding how a large language model predicts gendered pronouns. Specifically, given a sentence prefix including a traditionally gendered name (for example, “Mike” is often associated with males and “Jane” is often associated with females), how does the model predict what kind of pronoun should come after the sentence prefix? We will later see that understanding the mechanisms driving the model’s behavior on this benign task will yield insights for understanding gender-biased behavior of the model. Additionally, this investigation also provides an opportunity to test the ability of ObProp to accurately predict model behavior.
The gendered pronoun prediction problem was previously considered by Mathwin et al. (2023), where the authors used the “Automated Circuit Discovery” tool presented by Conmy et al. (2023) to investigate the flow of information between different components of GPT-2-small (Radford et al., 2019) in predicting subject pronouns (i.e. “he”, “she”, etc). We extend the problem setting in various ways. We investigate both the subject pronoun case (in which the model is to predict the token “she” versus “he”) and the object pronoun case (in which the model is to predict “her” versus “him”). Additionally, we seek to understand the underlying features responsible for this task, rather than just the model components involved, so that we can compare these features with the features that the model uses in producing gender-biased output.
@@
We investigate the model GPT-Neo-1.3B (Black et al., 2021), which has approximately 1.3B parameters, 24 layers, 16 attention heads, an embedding dimension of 2048, and an MLP hidden dimension of 8192. Note that ObProp is able to work with models that are significantly larger than those previously explored, such as GPT-2-small (117M parameters) (Radford et al., 2019), which has been the focus of recent interpretability work by Wang et al. (2022), inter alia.
@@
Table 1: The four attention heads with the greatest feature norms and path patching attributions (corrupted-clean logit differences) for both the $n_{\text{subj}}$ and $n_{\text{obj}}$ observables. $n_{\text{subj}}$ is the observable measuring the difference between the logits for " she" and " he"; $n_{\text{obj}}$ is the observable measuring the difference between the logits for " her" and " him". $l$::$k$ denotes the attention head with index $k$ at layer with index $l$. “Feature vector norms” refers to the norm of the feature vector associated with the attention head; “Path patching attributions” refers to the difference between the model’s output for the given observable when the given attention head’s activations was patched, and the model’s output for that given observable when the attention head was not patched.
@@ 4.2 Occupational gender bias
Now that we have understood some of the features relevant to predicting gendered pronous, we more directly consider the setting of occupational gender bias in language models, a widely-investigated problem (Bolukbasi et al., 2016; Vig et al., 2020). For a prompt like "My friend [NAME] is an excellent normal-…...", an LM which hasn’t been aligned using e.g. RLHF (Ouyang et al., 2022) is more likely to predict that the next token is " programmer" than " nurse" if [NAME] is replaced with a male name, and vice-versa for a female name (Brown et al., 2020). We applied observable propagation to the problem in order to go beyond prior work and understand the features responsible for this behavior. In particular, we considered the observable $n_{\text{bias}}=(e_{\text{{" nurse"}}}+e_{\text{{" teacher"}}}+e_{\text{{" secretary"}}})-(e_{\text{{" programmer"}}}+e_{\text{{" engineer"}}}+e_{\text{{" doctor"}}})$; this observable represents the extent to which the model predicts stereotypically-female occupations instead of stereotypically-male ones.
@@
More surprising is that the feature vector for $n_{\text{bias}}$ for 6::6$\to$9::1$\to$ had a cosine similarity of 0.966 with the feature vector for $n_{\text{subj}}$ for 6::6$\to$9::1$\to$13::11. Similarly, the $n_{\text{bias}}$ feature vector for mlp1$\to$6::6$\to$9::1$\to$ had a cosine similarity of 0.977 with the $n_{\text{subj}}$ feature vector for mlp1$\to$6::6$\to$9::1$\to$13::11. This indicates that the model uses the same features to identify both gender pronouns and likely occupations, given a traditionally-gendered name.
To determine the extent to which these feature vectors reflected model behavior, we ran the model on an artificial dataset of 600 prompts involving gendered names (see Appendix B), recorded the dot product of the model’s activations on the name token with the feature vectors, and recorded the model’s output with respect to the observables. The results can be found in Figure 2. Note that the correlation coefficient $r^{2}$ between the dot product with the $n_{\text{bias}}$ feature vector and the actual model output is 0.88, indicating that the feature vector is a very good predictor of model output.
@@
We then investigated the tokens in a 1M-token subset of The Pile that maximally activated the $n_{\text{bias}}$ feature vector. These tokens were primarily female names: tokens like " Rita", " Catherine", and " Mary", along with female name suffixes like "a" (as in “Phillipa”), "ine" (as in “Josephine”), and "ia" (as in “Antonia”). Surprisingly, the least-activating tokens were generally male common nouns, such as " husband", " brother", and " son" – but also words like " his", and even " male". This evidence even further supports the hypothesis that the model specifically uses gendered features in order to determine which occupations are most likely to be associated with a name. However, it is worth noting that part of the power of ObProp is that it allows us to test hypotheses such as this without needing to run the model on large datasets and record the tokens with the highest feature vector activations: simply by virtue of the extremely high cosine similarity between the $n_{\text{subj}}$ feature vector and the $n_{\text{bias}}$ feature vector, we could infer that the model was using gendered information to predict occupations. As such, looking at the maximally-activating tokens primarily served as a “sanity check”, verifying that the feature vectors returned by ObProp are human-interpretable.
@@ 4.3 Quantitative analysis across observables
We now evaluate ObProp’s performance across a broader variety of tasks, including subject pronoun prediction, identifying American politicians’ party affiliations, and distinguishing between C and Python code. We compare the performance of these feature vectors to the performance of feature vectors obtained by linear/logistic regression, a more traditional method (used by e.g. Kim et al. (2018) and others), but a more data-intensive one. For the former two tasks, we evaluate the correlation between the feature vectors and model outputs on the aforementioned artificial dataset used in the subject pronoun prediction experiments, and on an artificial dataset comprised of 40 Democratic politicians and 40 Republican politicians. For the programming language classification task, we evaluate the effectiveness of feature vectors in differentiating C and Python code using a natural dataset and the “Area Under the ROC Curve” metric.
The results are given in Table 3. For the subject pronoun prediction task, in order for the feature vector found by linear regression to match the performance of the ObProp feature vector, 60 prompts’ worth of embeddings had to be used for training; similarly, for the C vs. Python classification task, the logistic regression had to be trained on 50 code snippets’ worth of embeddings to obtain equal performance. In the political party prediction task, even when training on 3/4 of the dataset, the linear regression feature vector’s performance on the test set was well below that of the ObProp feature vector’s performance on the whole dataset. This suggests the ability of ObProp to match the performance of prior methods for finding feature vectors, and outcompete them especially in the low-data regime.
Task ObProp Logistic regression (trained) Regression training set size
Subject pronoun prediction $r^{2}\approx 0.945$ $r^{2}\approx 0.945$ 60 prompts
Political party prediction $r^{2}\approx 0.427$ $r^{2}\approx 0.295$ 60 prompts (3/4 of dataset)
C vs. Python AUC $\approx 0.9974$ AUC $\approx 0.9971$ 50 code snippets
@@ 5 Conclusion and Discussion
This culminated in a demonstration that the model specifically uses the feature of “gender” to predict the occupation associated with a name. Notably, even though experiments on larger datasets further supported this claim, observable propagation alone was able to provide striking evidence of this using minimal amounts of data. We hope that our approach, being independent of data, can democratize interpretability research and facilitate broader-scale investigations.
Furthermore, the conclusion that the model uses the same mechanisms to predict grammatical gender as it does to predict occupations portends difficulties in attempting to “debias” the model. This means that inexpensive inference-time attempts to remove bias from the model will likely also decrease model performance on desired tasks like correct gendered pronoun prediction (see Appendix F for additional experiments.) This reveals a clear future work direction to invest in more powerful methods, to ensure that models are both unbiased and useful.
Note that although ObProp demonstrates significant promise in cheaply unlocking the internal computations of language models, it does have limitations. In particular, ObProp only addresses the OV circuits of Transformers, ignoring computations in QK circuits responsible for mechanisms such as “induction heads” (Elhage et al., 2021). However, even though QK circuits are responsible for moving information around in Transformers, OV circuits are where computation on this information occurs. Thus, whenever we want to understand what sort of information the model uses to predict one token as opposed to another, the answer to this question lies in the model’s OV circuits, and ObProp can provide such answers. Given the power that the current formulation of ObProp has demonstrated already in our experiments, we are very excited about the potential for this method, and methods building upon it, to yield even greater insights in the near future.
Ethics statement
@@
A proof of Theorem 1 is given in Appendix H; a proof of Theorem 2 is given in Appendix J. Details on the datasets that we used in our experiments can be found in Appendix B. Further details regarding the experiments in Section 4.3 can be found in Appendix G. Details on how we chose the $x_{0}$ point used to approximate nonlinearities (as described in §2.3) can be found in Appendix D; for LayerNorm linear approximations, we used the estimation method described in Appendix C.1. Code for experiments can be found at github.com/jacobdunefsky/ObservablePropagation.
Appendix A Formal definition of observables
For further clarification, in this section, we provide a more formal definition of an observable.
Definition A.1.
An observable is a linear functional $n:\mathbb{R}^{\mathtt{d\_vocab}}\to\mathbb{R}$, where d_vocab is the number of tokens in the model’s vocabulary. We refer to the action of taking the inner product of the model’s output logits with an observable $n$ as getting the output of the model under the observable $n$.
Note that because all observables are linear functionals on a finite vector space, they can be written as row vectors. As such, it is often convenient to abuse notation, and associate an observable with its corresponding vector.
Observables often correspond to tasks on which we want to measure model output. The following example demonstrates how observables corresponding to a specific task can be constructed.
Example A.1.
Consider the task of predicting gendered subject pronouns. We want to measure the difference between the model’s predicted logits for the pronoun “she” and the pronoun “he”; this corresponds to how much more likely the model thinks that the next token will be the female subject pronoun over the male subject pronoun. Then if $e_{\mathtt{token}}$ is the one-hot vector with size d_vocab and a one in the position corresponding to token, then an observable corresponding to this task is given by $n=e_{\mathtt{"she"}}-e_{\mathtt{"he"}}$. This is because the output of the model under this observable precisely corresponds to the desired logit difference.
Appendix B Datasets
@@
  • Prompt templates for $n_{\text{subj}}$ (inspired by Mathwin et al. (2023)):
@@
A dataset of prompts was then generated by replacing the [NAME] substring in each prompt template with a name from a set of traditionally-male names and a set of traditionally-female names. These names were obtained from the “Gender by Name” dataset from UCI Machine Learning Repository (2020), which provided a list of names, the gender traditionally associated with each name, and a measure of the frequency of each name. The top 100 single-token traditionally-male names and top 100 single-token traditionally-female names from this dataset were collected; this comprised the list of names that we used.
Appendix C More on LayerNorm gradients
C.1 LayerNorm gradients are inversely proportional to input norms

### v2

@@ 4 Experiments
Armed with our ?observable propagation? toolkit for obtaining and analyzing feature vectors, we now turn our attention to the problem of gender bias in LLMs in order to determine the extent to which these tools can be used to diagnose the causes of unwanted behavior.
@@ 4.1 Gendered Pronouns Prediction
We first consider the related question of understanding how a large language model predicts gendered pronouns. Specifically, given a sentence prefix including a traditionally=gendered name (for example, ?Mike? is often associated with males and ?Jane? is often associated with females), how does the model predict what kind of pronoun should come after the sentence prefix? We will later see that understanding the mechanisms driving the model’s behavior on this benign task will yield insights for understanding gender-biased behavior of the model. Additionally, this investigation also provides an opportunity to test the ability of ObProp to accurately predict model behavior.
The gendered pronoun prediction problem was previously considered by Mathwin et al. (2023), where the authors used the ?Automated Circuit Discovery? tool presented by Conmy et al. (2023) to investigate the flow of information between different components of GPT-2-small (Radford et al., 2019) in predicting subject pronouns (i.e. ?he?, ?she?, etc). We extend the problem setting in various ways. We investigate both the subject pronoun case (in which the model is to predict the token ?she? versus ?he?) and the object pronoun case (in which the model is to predict ?her? versus ?him?). Additionally, we seek to understand the underlying features responsible for this task, rather than just the model components involved, so that we can compare these features with the features that the model uses in producing gender-biased output.
@@
We investigate the model GPT-Neo-1.3B (Black et al., 2021), which has approximately 1.3B parameters, 24 layers, 16 attention heads, an embedding dimension of 2048, and an MLP hidden dimension of 8192. Note that ObProp is able to work with models that are significantly larger than those previously explored, such as GPT-2-small (117M parameters) (Radford et al., 2019), which has been the focus of recent interpretability work by (Wang et al., 2022), inter alia.
@@
Table 1: The four attention heads with the greatest feature norms and path patching attributions (corrupted-clean logit differences) for both the $n_{\text{subj}}$ and $n_{\text{obj}}$ observables. $n_{\text{subj}}$ is the observable measuring the difference between the logits for " she" and " he"; $n_{\text{obj}}$ is the observable measuring the difference between the logits for " her" and " him". ?$l$::$k$? denotes the attention head with index $k$ at layer with index $l$. ?Feature vector norms? refers to the norm of the feature vector associated with the attention head; ?Path patching attributions? refers to the difference between the model’s output for the given observable when the given attention head’s activations was patched, and the model’s output for that given observable when the attention head was not patched.
@@ 4.2 Occupational Gender Bias
Now that we have understood some of the features relevant to predicting gendered pronous, we more directly consider the setting of occupational gender bias in language models, a widely-investigated problem (Bolukbasi et al., 2016; Vig et al., 2020). For a prompt like "My friend [NAME] is an excellent ...", an LM which hasn’t been aligned using e.g. RLHF (Ouyang et al., 2022) is more likely to predict that the next token is " programmer" than " nurse" if [NAME] is replaced with a male name, and vice-versa for a female name (Brown et al., 2020). We applied observable propagation to the problem in order to go beyond prior work and understand the features responsible for this behavior. In particular, we considered the observable $n_{\text{bias}}=(e_{\text{{" nurse"}}}+e_{\text{{" teacher"}}}+e_{\text{{" secretary"}}})-(e_{\text{{" programmer"}}}+e_{\text{{" engineer"}}}+e_{\text{{" doctor"}}})$; this observable represents the extent to which the model predicts stereotypically-female occupations instead of stereotypically-male ones.
@@
More surprising is that the feature vector for $n_{\text{bias}}$ for 6::6$\to$9::1$\to$ had a cosine similarity of 0.966 with the feature vector for $n_{\text{subj}}$ for 6::6$\to$9::1$\to$13::11. Similarly, the $n_{\text{bias}}$ feature vector for mlp1$\to$6::6$\to$9::1$\to$ had a cosine similarity of 0.977 with the $n_{\text{subj}}$ feature vector for mlp1$\to$6::6$\to$9::1$\to$13::11. This indicates that the model uses the same features to identify both gendered pronouns and likely occupations, given a traditionally-gendered name.
To determine the extent to which these feature vectors reflected model behavior, we ran the model on an artificial dataset of 600 prompts involving gendered names (see Appendix A), recorded the dot product of the model’s activations on the name token with the feature vectors, and recorded the model’s output with respect to the observables. The results can be found in Figure 1. Note that the correlation coefficient $r^{2}$ between the dot product with the $n_{\text{bias}}$ feature vector and the actual model output is 0.88, indicating that the feature vector is a very good predictor of model output.
@@
We then investigated the tokens in a 1M-token subset of The Pile that maximally activated the $n_{\text{bias}}$ feature vector [^2] If $y$ is a feature vector, then the maximally-activating tokens for $y$ are given by $\textrm{argmax}_{i}(y^{T}x_{i})$ where $x_{i}$ is the model’s hidden state on the $i$-th token in the dataset.. These tokens were primarily female names: tokens like " Rita", " Catherine", and " Mary", along with female name suffixes like "a" (as in ?Phillipa?), "ine" (as in ?Josephine?), and "ia" (as in ?Antonia?). Surprisingly, the least-activating tokens were generally male common nouns, such as " husband", " brother", and " son" – but also words like " his", and even " male". This evidence even further supports the hypothesis that the model specifically uses gendered features in order to determine which occupations are most likely to be associated with a name. However, it is worth noting that part of the power of ObProp is that it allows us to test hypotheses such as this without needing to run the model on large datasets and record the tokens with the highest feature vector activations: simply by virtue of the extremely high cosine similarity between the $n_{\text{subj}}$ feature vector and the $n_{\text{bias}}$ feature vector, we could infer that the model was using gendered information to predict occupations. As such, looking at the maximally-activating tokens primarily served as a ?sanity check?, verifying that the feature vectors returned by ObProp are human-interpretable.
@@ 4.3 Quantitative Analysis Across Observables
We now evaluate ObProp’s performance across a broader variety of tasks, including subject pronoun prediction, identifying American politicians’ party affiliations, and distinguishing between C and Python code. We use ObProp to find feature vectors for each task; for comparison, we also find feature vectors using the more data-intensive methods of linear/logistic regression and mean difference, standard methods used by Kim et al. (2018), Tigges et al. (2023) and many others. For the pronoun prediction task, we use the same artificial dataset used in the subject pronoun prediction experiments; for the political affiliation task, we use an artificial dataset comprised of 40 Democratic politicians’ and 40 Republican politicians’ names and consider the model’s logit difference between the tokens " Democrat" and " Republican". For the programming language classification task, we use a natural dataset of code. For each feature vector $y$, we look at the dot product $y^{T}x$ across inputs $x$. For the former two tasks, we evaluate the correlation between $y^{T}x$ and the model’s output; on the latter task, we apply the ?AUC-ROC? metric to evaluate the accuracy of $y^{T}x$ in differentiating between C and Python code.
The results are given in Table 3. For the subject pronoun prediction task, in order for the feature vector found by linear regression to match the performance of the ObProp feature vector, 60 prompts’ worth of embeddings had to be used for training; similarly, for the C vs. Python classification task, the logistic regression had to be trained on 50 code snippets’ worth of embeddings to obtain equal performance. In the political party prediction task, even when training on 3/4 of the dataset, the linear regression feature vector’s performance on the test set was well below that of the ObProp feature vector’s performance on the whole dataset. This suggests the ability of ObProp to match the performance of prior methods for finding feature vectors, and outcompete them especially in the low-data regime. For more details on these experiments, refer to Appendix B.
Task ObProp Logistic regression (trained) Mean difference vector (trained) Training set size
Subject pronouns $r^{2}\approx 0.945$ $r^{2}\approx 0.945$ $r^{2}\approx 0.899$ 60 prompts
Political parties $r^{2}\approx 0.427$ $r^{2}\approx 0.295$ $r^{2}\approx 0.0605$ 60 prompts (3/4 of dataset)
C vs. Python AUC $\approx 0.9974$ AUC $\approx 0.9971$ AUC $\approx 0.9052$ 50 code snippets
@@ 5 Conclusion and Discussion
This culminated in a demonstration that the model specifically uses the feature of ?gender? to predict the occupation associated with a name. Notably, even though experiments on larger datasets further supported this claim, observable propagation alone was able to provide striking evidence of this using minimal amounts of data. We hope that our approach, being independent of data, can democratize interpretability research and facilitate broader-scale investigations.
Furthermore, the conclusion that the model uses the same mechanisms to predict grammatical gender as it does to predict occupations portends difficulties in attempting to ?debias? the model. This means that inexpensive inference-time attempts to remove bias from the model will likely also decrease model performance on desired tasks like correct gendered pronoun prediction (see Appendix G for additional experiments.) This reveals a clear future work direction to invest in more powerful methods, to ensure that models are both unbiased and useful.
Note that although ObProp demonstrates significant promise in cheaply unlocking the internal computations of language models, it does have limitations. In particular, ObProp currently primarily addresses the OV circuits of Transformers, ignoring computations in QK circuits responsible for mechanisms such as ?induction heads? (Elhage et al., 2021). However, even though QK circuits are responsible for moving information around in Transformers, OV circuits are where computation on this information occurs. Thus, whenever we want to understand what sort of information the model uses to predict one token as opposed to another, the answer to this question lies in the model’s OV circuits, and ObProp can provide such answers.
Given the power that the current formulation of ObProp has demonstrated already in our experiments, we are very excited about the potential for this method, and methods building upon it, to yield even greater insights in the near future.
A note on SAEs
Sparse autoencoders (SAEs), as described by Cunningham et al. (2023) and Bricken et al. (2023), have recently made waves in the mechanistic interpretability community. SAEs can be trained unsupervised on model hidden states in order to yield a set of feature vectors that can represent any hidden state as a sparse linear combination of these feature vectors. Although we do find SAEs to be very promising, we fear that SAEs might incur some philosophical risks by assuming the existence of a set of ?ground-truth features? that can be found. Plausibly, SAEs might fail to account for cases where later-layer features depend on a dense subset of earlier-layer features or where feature vectors are composed in unintuitive ways to compute a task. As such, ObProp takes a different approach that privileges fidelity to computation over data: ObProp aims to find feature vectors (approximately) corresponding to the computation of human-interpretable tasks, the behavior of which feature vectors can then be quantifiably understood. In this manner, we hope to provide a complementary approach to the SAE paradigm (and one that could perhaps be integrated with it) in order to account for potential shortcomings of the latter.
Supplementary Statements
Impact Statement
@@
A proof of Theorem 1 is given in Appendix E.5; a proof of Theorem 2 is given in Appendix H. Details on the datasets that we used in our experiments can be found in Appendix A. Further details regarding the experiments in Section 4.3 can be found in Appendix B. Details on how we chose the $x_{0}$ point used to approximate nonlinearities (as described in §2.3) can be found in Appendix C; for LayerNorm linear approximations, we used the estimation method described in Appendix E.3. Code is available at https://github.com/jacobdunefsky/ObservablePropagation.
Appendix A Datasets
@@
  • Prompt templates for $n_{\text{subj}}$ (inspired by (Mathwin et al., 2023)):
@@
A dataset of prompts was then generated by replacing the [NAME] substring in each prompt template with a name from a set of traditionally-male names and a set of traditionally-female names. These names were obtained from the ?Gender by Name? dataset from (UCI Machine Learning Repository, 2020), which provided a list of names, the gender traditionally associated with each name, and a measure of the frequency of each name. The top 100 single-token traditionally-male names and top 100 single-token traditionally-female names from this dataset were collected; this comprised the list of names that we used.
Appendix B Experimental Details for Section 4.3
B.1 Datasets
The dataset used in the subject pronoun prediction task is the same artificial dataset described in Appendix A.
The dataset used in the American political party prediction task is an artificial dataset consisting of prompts of the form "[NAME] is a", where [NAME] is replaced by the name of a politician drawn from a list of 40 Democratic Party politicians and 40 Republican Party politicians. These politicians were chosen according to the list of ?the most famous Democrats? and ?the most famous Republicans? for Q3 2023 compiled by YouGov, available at https://today.yougov.com/ratings/politics/fame/Democrats/all and https://today.yougov.com/ratings/politics/fame/Democrats/all. The intuition behind this choice of dataset is that the model would be more likely to identify the political affiliation of well-known politicians, because better-known politicians would be more likely to occur in its training data. This is the primary reason that a smaller dataset is being used.
B.2 Task Definition
B.3 Feature Vectors
The ?mean difference? feature vectors for each task were calculated as follows. First, run the model on inputs from one class (e.g. female names, Democratic politicians, C code) and compute the mean vector across these inputs of the model embeddings at the same layer as the ObProp feature vector. Then, run the model on inputs from the other class (e.g. male names, Republican politicians, Python code) and compute the mean vector. Now, the ?mean difference? feature vector is simply the difference between these two mean vectors.
B.4 Task Evaluation
Appendix C Details on Linear Approximations for MLPs
Thus, we have two points $x_{-}$ and $x_{+}$ such that $g(x_{-})<0$ and $g(x_{+})>0$. Since MLPs are continuous, there therefore must be some point $x_{0}$ on the line between $x_{-}$ and $x_{+}$ at which $g(x_{0})=0$; this point lies at the intersection of the line between $x_{-}$ and $x_{+}$ with the MLP’s decision boundary. It stands to reason that the gradient at this decision boundary is more likely to capture the larger-scale behavior of the MLP and is less likely to be saturated, when compared to the gradient at more ?extreme? points like $x_{-}$ and $x_{+}$. Such an $x_{0}$ can be found using constrained optimization methods; we use the Python library MDMM (Crowson, 2021) to do so.
This approach given here is used in Line 14 of Algorithm 1 for dealing with MLP nonlinearities. For LayerNorms, we simply take the gradient at an input point $x_{-}$ or $x_{+}$.
Appendix D Details on Path Patching for Finding Important Computational Paths
In Section 4, we use path patching (Goldowsky-Dill et al., 2023) to determine which computational paths in the model are most important for a given task. Here, we provide more details on our implementation of path patching to measure the importance of a single computational path, along with more details on how we used path patching to find a set of important computational paths.
D.1 Path patching
Path patching is a causal method for determining the importance of a computational path in the model according to a given metric. This metric can generally be any function of the model’s outputs (e.g. the metric could be the cross-entropy next-token-prediction loss), but in this paper, our metrics are the dot product of the model’s logits with a given observable. We thus use path patching for determining how important a computational path is to the observable corresponding to a given task.
The high-level approach to path patching (which is common to other causal methods) is as follows. We consider two inputs to the model, a clean input and a dirty input, which display different behavior with respect to the metric. (For instance, if the metric is the dot product of the model’s logits with the $n_{\text{subj}}$ observable, then the clean input would be a prompt containing a traditionally-male name, and the dirty input would be a prompt containing a traditionally-female name.) First, run the model on both the clean and the dirty input, storing the model’s hidden states on both inputs. Then, take the hidden states from the clean run, replace certain hidden states with the corresponding ones from the dirty run, and re-run the model on this modified set of hidden states. After doing so, measure the difference, with respect to the given metric, between the model’s output on the clean input and the model’s output using the modified hidden states.
Our implementation of path patching is slightly different from the implementation described in Goldowsky-Dill et al. (2023): the implementation described here avoids the costly ?Treeify? function from the original.
First, we will describe how to perform path patching for a single-edge path. Given an earlier-layer component [^3] A component is an attention head or an MLP sublayer $c$ and a later-layer component $c^{\prime}$, let $o_{\text{clean}}$ be the output of $c$ on the clean input, and let $x^{\prime}_{\text{clean}}$ be the hidden states on the clean input before $c^{\prime}$. Similarly, let $o_{\text{dirty}}$ be the output of $c$ on the dirty input. Now, as explained by Elhage et al. (2021), a transformer’s hidden state before any component can be decomposed as the sum of all previous components’ outputs (along with the original token embedding and positional embedding). Thus, to measure the direct effect of the computational edge from $c$ to $c^{\prime}$, we replace $x^{\prime}_{\text{clean}}$ with $x^{\prime}_{\text{clean}}-o_{\text{clean}}+o_{\text{dirty}}$. This corresponds to replacing the contribution of $c$ to the input of $c^{\prime}$ with the output of $c$ on the dirty input. The path patching score is then computed by running the model with this modified hidden state and measuring the difference between its output and the clean output.
Now, this can be extended to longer computational paths as follows. Given a computational path of components $c^{(1)},\dots,c^{(k)}$, let $x^{(i)}_{\text{clean}}$ be the hidden states on the clean input before $c^{(i)}$, let $o^{(i)}_{\text{clean}}$ be the output of $c^{(i)}$ on the clean input, and let $o^{(i)}_{\text{dirty}}$ be the output of $c^{(i)}$ on the dirty input. Just as before, replace $x^{(2)}_{\text{clean}}$ with $x^{(2)}_{\text{dirty}}=x^{(2)}_{\text{clean}}-o^{(1)}_{\text{clean}}+o^{(1)}_{\text{dirty}}$. Run $c^{(2)}$ on $x^{(2)}_{\text{dirty}}$ and store the output as $o^{(2)}_{\text{dirty}}$. Repeat this process for the later components: replace $x^{(i)}_{\text{clean}}$ with $x^{(i)}_{\text{dirty}}=x^{(i)}_{\text{clean}}-o^{(i-1)}_{\text{clean}}+o^{(i-1)}_{\text{dirty}}$, run $c^{(i)}$ on $x^{(i)}_{\text{dirty}}$, and store the output as $o^{(i)}_{\text{dirty}}$. The path patching score is then given by the difference in the model’s output on these dirty hidden states and the model’s original output on the clean input.
D.2 Finding Important Computational Paths
Now that we can use path patching to allow us to determine the importance of a given computational path, we use the following greedy method in conjunction with path patching to find a set of important computational paths.
First, use path patching to identify the $k$ most important single-edge computational paths. Then, for each of those paths, identify the $k$ most important paths with the current path as a suffix; this gives us $k^{2}$ total paths. Now, take the top $k$ paths from these $k^{2}$ total paths, and repeat for as many iterations as one needs (until one has paths of a desired length).
The complexity of this process is $O(nk(pm+m\log m))$, where $n$ is the number of iterations, $k$ is the number of paths, $m$ is the number of nodes in the full computational graph of the model, and $p$ is the cost of path patching for one computational path. (To see this: at each iteration, we do path patching on $k$ parent paths, which has cost $p$ for each computational node $m$ in the model. Then, we sort the $m$ nodes to get the top $k$ paths, which gives us the $m\log m$ term.)
Note that this process can be made more efficient by using faster alternatives to path patching such as edge attribution patching (Syed et al., 2023).
Appendix E More on LayerNorms
In this section, we put forth various results relevant for the discussion of LayerNorm gradients in §2.4.
E.1 LayerNorm Input Norms per Layer
We calculated the average norms of inputs to each LayerNorm sublayer in the model, over the activations obtained from the 600 subject pronoun prompts in the artificial dataset described in §A. The results can be found in Figure 2. The wide variation in the input norms across different layers implies that input norms must be taken into account in any approximation of LayerNorm gradients.
Figure 2: Mean norms of activations before each LayerNorm
E.2 LayerNorm Weight Values Are Very Similar
In §2.4, the LayerNorm nonlinearity is defined as $\operatorname{LayerNorm}(x)=\frac{x-(\vec{1}^{T}x)\vec{1}}{\left\|x-(\vec{1}^{T}x)\vec{1}\right\|}+b$, where $\vec{1}$ is the vector of all ones. However, in actual models, after every LayerNorm operation as defined above, the output is multiplied by a fixed scalar constant equal to $\sqrt{d}$ (where $d$ is the embedding diension), multiplied by a learned diagonal matrix $W$, and then added to a learned bias vector $b$. Thus, the actual operation implemented is $\sqrt{d}W\operatorname{LayerNorm}(x)+b$, where $W$ is the learned diagonal matrix and $b$ is the learned vector.
This is important with regard to our earlier discussion of the extent to which LayerNorm affects feature vector directions, because although $b$ does not affect the gradient, nevertheless, if the values of $W$ are different from one another, this could cause the gradient to point in a different direction from the original feature vector.
However, empirically, we find that most of the entries in $W$ are very close to one another. This suggests that we can approximate $W$ as a scalar, meaning that $W$ primarily scales the gradient, rather than changing its direction. Therefore, if we want to analyze the directions of feature vectors rather than their magnitudes, then we can largely do so without worrying about LayerNorms.
In particular, we found that the average variance of scaling matrix entries across all LayerNorms in GPT-Neo-1.3B is $0.007827$. To determine the extent to which this variance is large, we calculated the ratio of the variance of each LayerNorm’s weight matrix’s entries to the mean absolute value of each layer’s embeddings’ entries. The results can be found in Figure 3. Note that the highest value found was $0.0714$ at Layer 0 – meaning that the average entry in Layer 0 embeddings was over 14.01 times larger than the variance between entries in that layer’s ln_1 LayerNorm weight. This supports our assertion that LayerNorm scaling matrices can be largely treated as constants.
One possible guess as to why this behavior might be occurring is this: much of the computation taking place in the model does not occur with respect to basis directions in activation space. However, the diagonal LayerNorm weight matrices can only act on these very basis directions. Therefore, the weight matrices end up ?settling? on the same nearly-constant value in all entries.
E.3 LayerNorm Gradients Are Inversely Proportional to Input Norms